Economic: Please assist me with these. I am a bit confused about the information asked from the given data and table 1.1.5. What are the value i should be looking at? This is the most recent GDP NIPA table released as you can see. Are the questions asked in the quiz only about Q3 as the table itself doesn't show quarter 4 . (Note: the table NIPA 1.1.5 can be slide from left to right and from up to don. So here i just took the pictures of the first two information related to questions 1 & 2).

02B: Recent data on GDP co U.S. Bureau of Economic × | + ← → O 凸bea gov Table Table careq de198ste & sun= 1 8 1910:08 0:4&1 921=su vey&1903-5&1904=1975&1905-20178 1906sq&1911:0t eq de198 Add to the favorites bar by selecting or by importing your favorites. t Table 1.1.5. Gross Domestic Product Billions of dollars] Seasonally adjusted at annual rates Last Revised on: December 21, 2017 - Next Release Date January 26 2018 2016 2017 Line 4 Q3 Gross domestic product 8,287.2 18,325.2 18,538.0 18,729.1 18,905.5 19,057.7 19,250.0 19,500.6 2,494.9 12,571.5 12,755.0 12,899.4 13,056.9 13,191.6 13,307.0 13,429.1 4,060.74,046.9 4,108.5 4,134.4 4,195.9 4,230.8 4,247.2 4,301.4 1,383.41,382.5 1,401.1 1,420.2 1,440.2 1,443.2 1,456.6 1,477.6 2,677.3 2,664.3 2,707.4 2,714.2 2,755.7 2,787.6 2,790.6 2,823.8 8,434.2 8,524.6 8,646.5 8,765.0 8,861.0 8,960.7 9,059.8 9,127.7 3,067.7 3,031.6 3,023.1 3,048.0 3,126.2 3,128.7 3,178.1 3,249.2 2,992.32,989.4 3,010.9 3,031.5 3,056.7 3,128.9 3,173.3 3,207.3 2,319.72,291.2 2,311.2 2,329.1 2,333.7 2,383.4 2,433.6 2,468.4 504.3 504.6 508.7525.6 525.8 548.4 563.0 559.0 1,083.51,046.2 1,044.3 1,040.9 1,044.3 1,057.6 1,082.3 1,111.0 731.9740.4 758.2 762.5763.7777.4 788.2 798.4 672.6 698.3 699.7 702.4 723.0 745.5 739.7 738.9 2 Personal consumption expenditures 3 4 Goods Durable goods Nondurable goods Services 7 Gross private domestic investment 8 Fixed investment Nonresidential 10 Structures Equipment Intellectual property products 12 13 Residential 3:50 PM 0后무드] 1/20/2018 ^ 4
Answer #1

Question is asking answers for most recent data i.e. Year 2017 and Quarter 3 .

(a).Y = 19500.6 (b).C=13429.1 (c).I=3249.2

(d).G=3356.5 (e).X=2345.9 (f).M=2880.1

(g)=68.9% [(C/Y)*100] (h)=16.7% [(I/Y)*100]   (i)=17.2% [(G/Y)*100]

(j)=12.0% [(X/Y)*100]

Values are taken from Q3 2017 . These are the highlighted values in the righmost column.